OpenStudy (anonymous):

Factor this expression completely 9a4 - b2

OpenStudy (whpalmer4):

That's a difference of two squares; that's the result you get when you multiply something like \[(x-y)(x+y) = x^2 - xy + xy - y^2 = x^2 - y^2\] Can you figure out what that means for you in this problem?

OpenStudy (anonymous):

no will u explain it please sorry

OpenStudy (whpalmer4):

Okay, if we match up your equation \[9a^4 - b^2\]with mine \[x^2-y^2\]then \[x^2=9a^4\]and\[y^2 = b^2\]right?

OpenStudy (anonymous):

yes but I don't understand this at all

OpenStudy (anonymous):

im so sorry

OpenStudy (whpalmer4):

okay, can you solve my two little equations for x and y?

OpenStudy (anonymous):

sure wat r they

OpenStudy (whpalmer4):

\[x^2 = 9a^4\] \[y^2 = b^2\]

OpenStudy (anonymous):

okay that's my problem but I don't know how to solve it

OpenStudy (whpalmer4):

Well, all the second one says is that there is some expression, \(y\), that if you multiply it by itself, it gives you \(b^2\). What would that be?

OpenStudy (whpalmer4):

Isn't that just \(b\)? \(b*b = b^2\) right?

OpenStudy (anonymous):

o ya okay I get that

OpenStudy (whpalmer4):

Well, how about \[x^2 = 9a^4\]\(x\) is some expression that if you multiply \(x*x\) you get \(9a^4\). Any ideas? It will have a number, and it will have a letter...

OpenStudy (anonymous):

(x*x=x^2) (9a^4= ) im confused now

OpenStudy (whpalmer4):

okay, what is the square root of 9?

OpenStudy (anonymous):

I think 3 right @whpalmer4

OpenStudy (whpalmer4):

does 3*3= 9? if so, then the answer is "yes, I'm sure of it!" :-)

OpenStudy (anonymous):

okay how else would I do this

OpenStudy (whpalmer4):

Okay, so we're going to have 3<something>*3<something> = 9a^4 Any guesses about the <something>?

OpenStudy (anonymous):

3*3=9a^4

OpenStudy (whpalmer4):

3<something>*3<something> = 9<something>^2

OpenStudy (anonymous):

u confued me

OpenStudy (whpalmer4):

Do you remember the properties of exponents?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

9*9=9^2

OpenStudy (whpalmer4):

One of them is \[a^n*a^m = a^{n+m}\] Or if the two exponents are equal, that becomes \[a^n*a^n = a^{2n}\]

OpenStudy (whpalmer4):

So we want to multiply two equal things together and get \(a^4\)...what does that property I just showed you suggest we might use?

OpenStudy (anonymous):

distropit idk my to spell it

OpenStudy (whpalmer4):

no, I'm looking for a value...

OpenStudy (anonymous):

idk im confused

OpenStudy (whpalmer4):

Okay, \[a^n*a^n = a^{2n}\]We are trying to find something that when multiplied by itself gives you \(a^4\). Isn't \(a^4\) the same as \(a^{2n}\) if \(n=2\)?

OpenStudy (anonymous):

idk the letters r confusing me

OpenStudy (whpalmer4):

does 4 = 2n, if n = 2?

OpenStudy (anonymous):

can u draw it out

OpenStudy (whpalmer4):

Does 4 = 2*2?

OpenStudy (anonymous):

yes 2*2=4

OpenStudy (whpalmer4):

Okay, so does 4 = 2n, if n = 2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay we does that have to do with my math

OpenStudy (whpalmer4):

Excellent! :-) So that means that \[a^2*a^2 = a^4\] And we can package that up with \[3*3 = 9\]and we get \[3a^2*3a^2 = 9a^4\] So we can factor our expression as \[(3a^2-b)(3a^2+b) = 9a^4-b^2\]

OpenStudy (whpalmer4):

Because\[(3a^2-b)(3a^2+b) = 9a^4 + 3a^2b-3a^2b-b^2 = 9a^4 -b^2\]